The perimeter of a rectangle is 120 feet. Use an equation to find the length and width of the rectangle if the length is twice the width.
Use P = 2W + 2L.

Answer by TimothyLamb(4379)

2L + 2w = 120
L = 2w
---
2L + 2w = 120
2(2w) + 2w = 120
4w + 2w = 120
6w = 120
---
w = 20 ft
L = 40 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php